# Fermi acceleration under Lorentz invariance violation

### v1

@@ 3.1 UHECR energy spectrum
Equation 3.16 can be numerically solved, as shown in Appendix A. The solution is not straightforward and requires one extra assumption that the energy spectrum can be described by a power law in energy $N=N_{0}E^{\lambda(E)}$. The relation between the escape time and average time between collisions defines the index of the power law of the energy spectrum. We take $\frac{1}{\alpha\tau_{scp}}=8/3$ to match the index of $-2$ as suggested by the UHECR data.
@@ 4.1 UHECR energy spectrum
$\frac{dN}{dE}=-\left[1+\frac{2(1+\delta_{n}E^{n})-n\delta_{n}E^{n}}{2(1+\delta_{n}E^{n})^{3/2}}\right]\frac{\sqrt{1+\delta_{n}E^{n}}N}{E}\textrm{,}$ (4.4)
@@
The introduction of LIV in Fermi’s first-order mechanism causes a suppression of the flux of particles for energies above the break, similar to that observed in measurements [34]. Figures 5 and 6 show a fit of the LIV parameter to the energy spectrum measured by the Pierre Auger Observatory. The fit has the unique intention of comparing the suppression caused by LIV in the first-order Fermi mechanism with the data. It is not our intention to explain the data or limit LIV coefficients. The resulting spectrum of the fit could be seen as the flux of particles produced by one source near Earth. The fit resulted in $\delta_{1}\leq(5.1\pm 1.0)\times 10^{-20}\ \textrm{eV}^{-1}$ and $\delta_{2}\leq(2.9\pm 0.2)\times 10^{-39}\ \textrm{eV}^{-2}$.
For this system, we can write the acceleration time following the same definition as in equation 3.20
@@
Note that for the first order mechanism, the acceleration time will now increase as we go to higher energies. From this, we can extract the maximum energy
@@ 5 Conclusion
In this paper, we have calculates the acceleration of particles using the Fermi mechanisms of first and second order under the assumption of Lorentz invariance violation. LIV is introduced in a phenomenological approach by a modified dispersion relation. We calculated the resulting energy spectrum of particles accelerated in the source and the acceleration time.
For the second-order Fermi mechanism, the changes introduced by LIV entail a change in the spectral index for energies above the break. Such deviations depend on the order of the violation considered in equation 2.3. For the first-order mechanism, the changes introduced by LIV are important as the flux of particles is strongly suppressed for energies above the LI break, leading to a possible measurable effect. The suppression caused by LIV is of the same order of magnitude of the suppression measured in the data [34]. Therefore, LIV in the acceleration mechanism could also be important to explain the suppression seen in the data, adding a third hypothesis to the most common ones in the literature relating the suppression to the interaction with the photon background [37] or to the maximum power of the sources [36, 38].
@@
The calculation presented here can be used to derive limits on the LIV coefficients by using the measured UHECR energy spectrum. However, the calculation of the limits on the LIV coefficients depends on astrophysical assumptions such as the distribution of sources and the mass composition [14, 39]. The overall effect of LIV in the analysis of UHECR must include the three regions mentioned in the introduction: source, extragalactic medium and shower in the Earth’s atmosphere. There is no way that LI is broken only in one region and not in the others. With the calculations presented here, the fundamental knowledge to calculate the effect of LIV in UHECR is set.
@@ Appendix A Appendix
Figure 5: Energy spectrum of particles accelerated by first-order Fermi mechanisms with Lorentz Invariance Violation given by the $\delta$ parameter for the first $\delta_{1}$ order of the energy dispersion relation expansion. The blue dots show the energy spectrum measured by the Pierre Auger Observatory [34]. The $\delta_{1}$ parameter was fitted to the data.
Figure 6: Energy spectrum of particles accelerated by first-order Fermi mechanisms with Lorentz Invariance Violation given by the $\delta$ parameter for the first $\delta_{2}$ order of the energy dispersion relation expansion. The blue dots show the energy spectrum measured by the Pierre Auger Observatory [34]. The $\delta_{2}$ parameter was fitted to the data.

### v2

@@ 3.1 UHECR energy spectrum
Equation 3.16 can be numerically solved, as shown in Appendix A. The solution is not straightforward and requires one extra assumption that the energy spectrum can be described by a power law in energy $N=N_{0}E^{\lambda(E)}$. The relation between the escape time and average time between collisions defines the index of the power law of the energy spectrum. We take $\frac{1}{\alpha\tau_{scp}}=8/3$ to match the index of $-2$ as suggested by the UHECR data [31, 32].
@@ 4.1 UHECR energy spectrum
$\frac{dN}{dE}=-\left[1+\frac{2(1+\delta_{n}E^{n})-n\delta_{n}E^{n}}{2(1+\delta_{n}E^{n})^{3/2}}\right]\frac{\sqrt{1+\delta_{n}E^{n}}N}{E}\textrm{.}$ (4.4)
@@
The introduction of LIV in Fermi’s first-order mechanism causes a suppression of the flux of particles for energies above the break, similar to that observed in measurements done by the Pierre Auger Observatory [36]. Figures 5 and 6 show the energy spectrum measured by the Pierre Auger Observatory and the solution of equation 4.4 for different values of $\delta_{1}$ and $\delta_{2}$, respectvely. The plot is an illustration of the effect of LIV suppression in the Fermi first-order mechanism. It is not our intention to explain the data or limit LIV coefficients in this analysis.
The maximum energy that an acceleration region can provide is of utmost importance; therefore, the effects of LIV must also be considered. For this system, we can write the acceleration time following the same definition as in equation 3.20
@@
Note that for the first-order mechanism, the acceleration time will now increase as we go to higher energies. From this, we can extract the maximum energy
@@ 5 Conclusion
In this paper, we have calculated the acceleration of particles using the Fermi mechanisms of first and second order under the assumption of Lorentz invariance violation. LIV is introduced in a phenomenological approach by a modified dispersion relation. We calculated the resulting energy spectrum of particles accelerated in the source and the acceleration time.
For the second-order Fermi mechanism, the changes introduced by LIV entail a change in the spectral index for energies above the break. Such deviations depend on the order of the violation considered in equation 2.3. For the first-order mechanism, the changes introduced by LIV are important as the flux of particles is strongly suppressed for energies above the LI break, leading to a possible measurable effect. The suppression caused by LIV is of the same order of magnitude of the suppression measured in the data [36]. The suppression is typically explained by the interaction of particles with background photons [39] or by the maximum power of the sources [38, 40]. The addition of LIV in the acceleration mechanism could help complement these hypotheses without excluding them, as they are crucial effects at the highest energies.
@@
The calculation presented here can be used to derive limits on the LIV coefficients by using the measured UHECR energy spectrum. However, the calculation of the limits on the LIV coefficients depends on astrophysical assumptions such as the distribution of sources and the mass composition [14, 41]. The potential of UHECR data to find or impose limits on LIV will improved significantly if a source is identified and if a subset of particles is selected.
Most of the uncertainty in data analysis arises from the lack of knowledge regarding the type of particle arriving on Earth. The acceleration time, and consequently the maximum energy a particle can achieve at its source, is dependent on the particle’s charge. Additionally, the total distance traveled by the particle from the source to Earth is influenced by its charge due to deviations caused by magnetic fields.
Furthermore, the ongoing challenge in identifying individual ultra-high-energy cosmic ray (UHECR) sources introduces significant uncertainty. Strong assumptions about source distributions and their characteristics further exacerbate this issue. Identifying a point source would eliminate many unknowns. Even if only the class of objects (e.g., AGN or starburst regions) responsible for these emissions is identified, it would allow for a deeper investigation into the details of the acceleration mechanisms involved.
For these reasons, the planned upgrades to the Pierre Auger Observatory [42] and the Telescope Array Observatories [43] are expected to significantly enhance the detectability of Lorentz invariance violation (LIV) in the coming years. These upgrades aim to select a subset of proton events and identify local sources, thereby reducing the major uncertainties in the search for LIV.
The overall effect of LIV in the analysis of UHECRs must consider all three regions discussed in the introduction: the source, the extragalactic medium, and the atmospheric shower on Earth. It is implausible that Lorentz invariance is violated in only one of these regions and not the others. The calculations presented in this paper contribute to completing the overall scenario by incorporating considerations of the acceleration mechanism.
@@ Appendix A Appendix
Figure 5: Energy spectrum of particles accelerated by first-order Fermi mechanisms with Lorentz Invariance Violation given by the $\delta$ parameter for the first $\delta_{1}$ order of the energy dispersion relation expansion. The blue dots show the energy spectrum measured by the Pierre Auger Observatory [36]. The $\delta_{1}$ parameter was fitted to the data. Different values of $\delta_{1}$ are shown. All spectrums have a normalization factor of $N_{0}=\left(8.4\pm 0.3\right)\times 10^{41}$.
Figure 6: Energy spectrum of particles accelerated by first-order Fermi mechanisms with Lorentz Invariance Violation given by the $\delta$ parameter for the first $\delta_{2}$ order of the energy dispersion relation expansion. The blue dots show the energy spectrum measured by the Pierre Auger Observatory [36]. The $\delta_{2}$ parameter was fitted to the data. Different values of $\delta_{2}$ are shown. All spectrums have a normalization factor of $N_{0}=\left(7.16\pm 0.07\right)\times 10^{41}$.